Three pipes A,B and C in a tank. They can fill the tank in 10 hrs, 20 hrs and 25 hrs respectively.
At first all of them are opened simultaneously.
After 2 hrs,tap C is closed and A and B are kept running.
After 4 hrs, tap B is also closed. The remaining work is done by Tap A alone.
Find the percentage of the work done by Tap A by itself.

Found 2 solutions by Azeez majiyagbe, richwmiller:
Answer by Azeez majiyagbe(1)
Answer by richwmiller(17219)

2/10+2/20+2/25+4/10+4/20+x/10=1
x = 1/5 hour
.2/10=.02 or 2%
A did 2% of the work alone